# Li Tana,b , Shengrong Wanga,c

### v1

@@ 3 Existence and Uniqueness
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X_{s}^{n}|^{u_{2}}\bigg{)}\leq C+C\int_{0}^{2\tau}\left[\mathbb{E}\bigg{(}\sup_{s\in[0,r]}|X_{s-\tau}^{n}|^{u_{1}}\bigg{)}\right]^{\frac{(l+1)u_{2}}{u_{1}}}{\rm d}r\leq C.$
By induction, we can derive the assertion. ∎
Lemma 3.2.
Let Assumptions 2.2-2.5 hold. Then for $0\leq s\leq t\leq T$ with $t-s<1$, and $p>1/H$, $(l+1)p\leq\overline{p}$, $n\geq 2/\tau$, we have
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}\leq C[(t-s)^{p}+(t-s)^{pH}],$
where $C$ is a positive constant depending on $p,H,\xi,T,L$ but independent of $s$ and $t$.
Proof.
By (3.1), we may compute
$X_{t}^{n}-X_{s}^{n}=$ $D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})+\int_{s}^{t}b\left(X^{n}_{r-1/n},X^{n}_{r-\tau},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r+\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}.$
By the elementary inequality, the Hölder inequality, Assumptions 2.3, 2.4, 2.5 and using same techniques as (3.3), we get
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
$\leq$ $3^{p-1}\mathbb{E}|D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})|^{p}+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}b(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}r\bigg{\rvert}^{p}$
$+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{p}$
$\leq$ $3^{p-1}\bigg{[}\lambda^{p}\mathbb{E}|X_{t-\tau}^{n}-X_{s-\tau}^{n}|^{p}+(t-s)^{p-1}$
×$\times\mathbb{E}\int_{s}^{t}C^{p}\bigg{(}1+|X_{r-1/n}^{n}|^{l+1}+|X_{r-\tau}^{n}|^{l+1}+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}{\rm d}r$
$+C(t-s)^{pH-1}\mathbb{E}\int_{s}^{t}\bigg{(}1+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}.$
For $0\leq s\leq t\leq\tau$, by Assumption 2.2, we get
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
$\leq$ $3^{p-1}\lambda^{p}\mathbb{E}|\xi(t-\tau)-\xi(s-\tau)|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
$\leq$ $3^{p-1}\lambda^{p}K_{0}|t-s|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r.$
Since $(l+1)p\leq\overline{p}$, we can applying Lemma 3.1 to obtain the assertion. For $\tau\leq s\leq t\leq T$, noting that $01-3^{p-1}\lambda^{p}>$ 0, we get the desired assertion by applying Lemma 3.1 again. ∎
Theorem 3.3.
Let Assumption 2.4 hold. Then for $p>1/H$ and $(l+1)p\leq\overline{p}$, (2.5) admits a global unique solution.
Proof.
We split the proof into two steps.
Step 1. Existence. For $m>n\geq 2/\tau$ and $t\in[0,T]$, by (2.5), we may compute
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}$
$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}\sigma\big{(}\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-\sigma\big{(}\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{)}$
=:absent:=:= : $W_{1}+W_{2}+W_{3},$
By Assumption 2.3, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}\leq\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}.$
By the Hölder inequality, Assumption 2.4, we get
$W_{2}=$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
$\leq$ $(3t)^{p-1}\mathbb{E}\int_{0}^{t}\bigg{\lvert}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{\rvert}^{p}{\rm d}r$
$\leq$ $(3t)^{p-1}K_{3}\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}$
×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}+\mathbb{W}_{q}\big{(}\mathcal{L}_{X_{r-1/m}^{m}},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}^{p}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{p}$
×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{p}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
Since for $l>1$, we have $2pl\leq(l+1)p\leq\overline{p}$, then by the Hölder inequality, Lemmas 3.1, 3.2, we get
$W_{2}\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{2p}\bigg{]}^{1/2}$
×$\times\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{r-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r$
$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}.$
By Assumption 2.5, Lemmas 3.1, 3.2 and the same techniques as (3.3), we get
$W_{3}=$ $3^{p-1}\mathbb{E}\bigg{[}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{(}\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\bigg{)}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{]}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})|^{p}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}K_{4}^{p}(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/m}^{m}}-\mathcal{L}_{X_{r-1/n}^{n}}))^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{m}|^{p}{\rm d}r+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
Combining $W_{1},W_{2},W_{3}$, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u}^{m}-X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
Noting that $\lambda<1$, by the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
Define a sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)p2^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$. It is easy to see that $u_{i}\geq p$, $u_{[T/\tau]+1}=p$ and $2u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$. For $t\in[0,\tau]$, we get
(3.4) $\mathbb{E}\bigg{(}\sup_{s\in[0,\tau]}|X^{m}_{s}-X_{s}^{n}|^{u_{1}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
For $t\in[0,2\tau]$, by the Hölder inequality, we get
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}$
[…]
by (3.4),
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
By induction, we can see that for $t\in[0,T]$,
(3.5) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
Consequently, $(X^{n})_{n\geq 2/\tau}$ is a Cauchy sequence for $p>1/H$. Denote the limit of $(X^{n})$ by $X$ and letting $m\to\infty$ in (3.5),
(3.6) $\lim_{n\to\infty}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-X_{s}^{n}|^{p}\bigg{)}=0.$
Then, we should show that $X_{t}$ is a solution to (2.5). For $t\in[0,T]$, we get
$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}=$ $\mathbb{E}|X_{t}-X^{n}_{t}+X^{n}_{t}-X^{n}_{t-1/n}|^{p}$
$\leq$ $2^{p-1}\mathbb{E}|X_{t}-X^{n}_{t}|^{p}+2^{p-1}\mathbb{E}|X^{n}_{t}-X^{n}_{t-1/n}|^{p}.$
By Lemma 3.2 and (3.6), we get
$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}\to 0,\ \ {\rm as}\ \ n\to\infty.$
Taking $n\to\infty$ in (3.1), we get
$X_{t}-D(X_{t-\tau})=\xi(0)+D(\xi(-\tau))+\int_{0}^{t}b(X_{s},X_{s-\tau},\mathcal{L}_{X_{s}}){\rm d}s+\int_{0}^{t}\sigma(\mathcal{L}_{X_{s}}){\rm d}B_{s}^{H},\ \ t\in[0,T].$
Step 2. Uniqueness. Let $X_{t}$ and $Y_{t}$ be two solutions of (2.5) with $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$. By Lemma 2.7 again, we get
$|X_{t}-Y_{t}|^{p}\leq\lambda|X_{t-\tau}-Y_{t-\tau}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}-D(X_{t-\tau})+Y_{t}-D(Y_{t-\tau})|^{p},$
which implies,
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}$
$\leq$ $\sup_{-\tau\leq r\leq 0}|X_{r}-Y_{r}|^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}+\frac{1}{(1-\lambda)^{p-1}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}.$
Since $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$ and $\lambda\in(0,1)$,
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}\leq\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}=\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|\Pi(r)|^{p},$
where $\Pi(r)=X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})$. For $t\in[0,T]$, by the Itô formula and the elementary inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Pi(s)|^{p}\bigg{)}\leq$ $p\mathbb{E}\int_{0}^{t}|\Pi(r)|^{p-2}\langle\Pi(r),b(X_{r},X_{r-\tau},\mathcal{L}_{X_{r}})-b(Y_{r},Y_{r-\tau},\mathcal{L}_{Y_{r}})\rangle{\rm d}r$
$+p(p-1)H\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-2}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|^{2}{\rm d}r\bigg{)}$
$+p\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-1}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|{\rm d}B_{r}^{H}\bigg{)}.$
$=$ $J_{1}+J_{2}+J_{3}.$
By Assumption 2.4, the Young’s inequality, we get
$J_{1}=$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}|X_{r}-Y_{r}|^{2}+|X_{r-\tau}-Y_{r-\tau}|^{2}+(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
By Assumption 2.5 and the Young’s inequality, we get
$J_{2}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{[}(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
In the way in (3.3), by Assumption 2.5, the Young’s inequality, the Hölder inequality, we get
$J_{3}\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
Combining $J_{1}$-$J_{3}$ and applying the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-Y_{s}|^{p}\bigg{)}\leq C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
By induction, we derive that $X_{t}=Y_{t},t\in[0,T]$, $\mathbb{P}$-a.s. The proof is complete. ∎

### v2

@@ 3 Existence and Uniqueness
$\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right)+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r$
$+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
Choosing $\epsilon$ small enough so that $C\epsilon<\frac{1}{2}$ and absorbing the first term, it implies
$\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r$
$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
Since singular kernels $\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r$ are typically “stronger” than regular kernels $\int_{0}^{t}(t-r)^{0}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r$, we derive from (3.5) that
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)$
$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
Using the Gronwall inequality leads to
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)\leq Ce^{Ct}\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
By induction, we arrive at
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)\leq\frac{(Ce^{Ct}\Gamma(2H)t^{2H})^{k}}{\Gamma(2kH+1)}\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(1),(0)}|^{\overline{p}}\bigg).$
With the properties of the Gamma function and the fact that $\mathbb{E}\bigg(\sup\limits_{0\leq s\leq t}|\Delta_{s}^{(1),(0)}|^{\overline{p}}\bigg)\leq C$ derived by (3.1), we can show that $\{X_{t}^{(k)}\}$ is a Cauchy sequence, then the limit, denoted by $X$, is a solution of (2.5).
Step 3. Uniqueness Let $X_{t}$ and $Y_{t}$ be two solutions of (2.5) with $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$. By Lemma 2.7 again, we establish
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{\overline{p}}\leq\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq r\leq t}|\Upsilon_{r}|^{\overline{p}},$
where $\Upsilon_{r}=X_{r}-D(X_{r-\tau})-Y_{r}+D(Y_{r-\tau})$. Similar to the process of Step 2, it holds that
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}-Y_{s}|^{\overline{p}}\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}-Y_{s}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}-Y_{s}|^{\overline{p}}\bigg){\rm d}r.$
Further, we derive
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}-Y_{s}|^{\overline{p}}\bigg)\leq C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}-Y_{s}|^{\overline{p}}\bigg){\rm d}r.$
Although the kernel $(t-r)^{2H-1}$ is singular as 2H1>02H-1>0, it is integrable on [0, T]. Therefore, applying Lemma 2.8 with $a=0$, $\beta=2H$ and $\gamma=1$, we find that the generalized Mittag-Leffler function is bounded, guaranteeing a unique zero solution for the inequality. That is,
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}-Y_{s}|^{\overline{p}}\bigg)=0.$
Therefore, $X_{t}=Y_{t}$ almost surely for all $t\in[0,T]$. This completes the proof. ∎
Remark 3.2.
For the existence and uniqueness of the McKean-Vlasov NSDDEs, we adopt the Picard approximation method to construct a Cauchy sequence and show the existence and uniqueness of the solution. Since the super-linearly growing coefficients in our proposed model introduces additional technical complexities that demand refined analytical tools, our analysis differs fundamentally from those of related works. For McKean-Vlasov SDEs driven by standard Brownian motions (e.g., [17]) or by fractional Brownian motions without super-linearly growing coefficients (e.g., [18, 19]), the uniqueness is typically proven via direct estimates of the difference between two candidate solutions. In contrast, our model involves fractional noise and super-linear coefficients, which necessitates the use of the fractional Itô formula to properly account for the non-Markovian and rough nature of the fractional Brownian motions. This application of the fractional Itô formula not only resolves the technical hurdles posed by the fractional noise but also provides a generalizable framework for analyzing McKean-Vlasov SDEs driven by fractional Brownian motions.